# Symmetries and singular behaviors with Bohmian trajectories

### v1

@@ 3 Self-accelerating shape-invariant propagation
As it is well known, the free propagation of Gaussian wave packets in configuration space is characterized by a shape invariance of their amplitude, which remains scale invariant at any time if we take into account their hyperbolic widening. In 1979, Berry and Balazs [47] showed that it is also possible to find non-dispersive wave packet solutions to Schrödinger’s equation in configuration space. The only requirement to observe this propagation invariance is the initial amplitude of the wave packet to be described by an Airy function. The probability density of Airy wave packets presents two important symmetries: ($i$) shape invariance during propagation, unlike Gaussian wave packets, and ($ii$) self-acceleration without the presence of any external potential. Almost thirty years later, Siviloglou et al. [48] produced the first experimental evidence of this type of dispersionless wave packets with modulated light beams, making use of the isomorphism discussed in Sec. 2.2.
@@
One may wonder what happens with the above-mentioned symmetries is the beam is close to the shape of an Airy function, but not an Airy function strictly speaking. This may happen, for instance, if our imaging system has a limited extension and we cannot recreate with a high accuracy, on the input plane, the fast oscillations of the tail of the Airy beam as $x$ gets more and more negative. Let us thus consider that the beam tail is affected by an exponential decrease [48, 49], i.e., the initial field amplitude reads as
@@
where $y=x-z^{2}/4+i\gamma z$. It is by virtue of the argument $y$ that now the Airy function is a complex function and, therefore, apart from the linear contribution in $z$ to the velocity field, we are going to have an additional contribution arising from the phase developed with increasing $z$ by the Airy function. The effective velocity field thus acquire now the non-analytical functional form
@@ 5 Concluding remarks
We have also seen that, in order to determine the flow of probability or light intensity in space, a rather convenient tool is provided by the trajectories that arise after integrating along the propagation variable the above-mentioned velocity field. This renders the so-called Bohmian trajectories, which, again, have nothing to do (at least, in principle) with ontological hidden variables, but just with well-defined elements naturally arising within a hydrodynamic reformulation of Schrödinger and, in general, Schrödinger-type equations. These trajectories allow us to extract valuable local information about the propagation of the systems described by such equations, which, in turn, supply an additional perspective on how and why probabilities or light intensities redistribute spatially in they way the do it. This event-by-event picture of wave phenomena results worth emphasizing, because it enables the reconstruction of such phenomena under a single-event basis (single photons, single electrons, etc.), that is, in circumstances of a rather low flux of particles [51, 52, 53, 54], thus making clearer the connection between particle wave theories and statistics [55].
Here, in particular, we have considered these tools to study and explain the symmetries and singularities that characterize two types of challenging behaviors that arise in the simplest propagation conditions, i.e., under free propagation. Specifically, we have investigated those properties in two phenomena that have been recently considered in more detail elsewhere: ($i$) the lack of dispersion and self-acceleration of Airy beams during their propagation [25], and ($ii$) the spontaneous self-focusing displayed by a beam when a particular phase distribution is imprinted in it initially [26]. In the first case, the trajectories have allowed us to provide a quantitative measure to understand how Airy-type beams degrade when we introduce conditions that diminish their ideal initial shape. Thus, we find that, the translation symmetry along the hyperbolic Bohmian trajectories disappears gradually, because there is an effective transfer of flux towards the rear part of the beam, which eventually translates into an also loss of its transverse self-acceleration. In other words, what keeps the Airy beam properties is the presence of an infinite tail. This generates a constant pressure forwards [56, 57], which is evidenced by all trajectories, regardless of their initial position, bending over at a constant rate.

### v2

@@ 3 Self-accelerating shape-invariant propagation
A free-propagating Gaussian wave packet preserves its Gaussian amplitude in configuration space at any time, with its width undergoing a hyperbolic increase [9, 14]. Accordingly, for $t\gg 2m\sigma_{0}^{2}/\hbar$ (with $\sigma_{0}$ being its width at $t=0$), the amplitude essentially increases linearly with time, thus remaining scale invariant in spite of its progressive dispersion. In 1979, Berry and Balazs [47] showed that it is also possible to find non-dispersive wave packet solutions to Schrödinger’s equation in configuration space. The only requirement to observe this invariance property during their propagation is that the initial amplitude of the wave packet has to be described by an Airy function. The amplitude of these so-called Airy wave packets thus presents two important symmetries: ($i$) shape invariance during propagation, unlike Gaussian wave packets, and ($ii$) self-acceleration without the presence of any external potential. Almost thirty years later, Siviloglou et al. [48] produced the first experimental evidence of this type of dispersionless wave packets with modulated light beams, making use of the isomorphism discussed in Sec. 2.2.
@@
One may wonder what happens with the above-mentioned symmetries if the beam is close in shape to an Airy function, but it is not an Airy function strictly speaking. This can happen, for instance, when our imaging system has a limited extension and therefore it is not possible to recreate with a high accuracy, on the input plane, the fast oscillations of the tail of the Airy beam as $x$ gets more and more negative. Let us thus consider that the beam tail is affected by an exponential decrease [48, 49], i.e., the initial field amplitude reads as
@@
where $y=x-z^{2}/4+i\gamma z$. By virtue of the complex argument $y$, now the Airy function also becomes complex valued. Therefore, the velocity field will consist of the linear $z$-dependent contribution plus an additional term arising from the phase developed by the Airy function with increasing $z$. More specifically, this effective velocity field acquires the non-analytical functional form
@@ 5 Concluding remarks
We have also seen that, in order to determine the flow of probability or light intensity in space, a rather convenient tool is provided by the trajectories that arise after integrating along the propagation variable the above-mentioned velocity field. This renders the so-called Bohmian trajectories, which, again, have nothing to do (at least, in principle) with ontological hidden variables, but just with well-defined elements naturally arising within a hydrodynamic reformulation of Schrödinger and, in general, Schrödinger-type equations. These trajectories allow us to extract valuable local information about the propagation of the systems described by such equations, which, in turn, supply an additional perspective on how and why probabilities or light intensities redistribute spatially in the way they do it. This event-by-event picture of wave phenomena results worth emphasizing, because it enables the reconstruction of such phenomena under a single-event basis (single photons, single electrons, etc.), that is, in circumstances of a rather low flux of particles [51, 52, 53, 54], thus making clearer the connection between particle wave theories and statistics [55].
Here, in particular, we have considered these tools to study and explain the symmetries and singularities that characterize two types of challenging behaviors that arise in the simplest propagation conditions, i.e., under free propagation. Specifically, we have investigated those properties in two phenomena that have been recently considered in more detail elsewhere: ($i$) the lack of dispersion and self-acceleration of Airy beams during their propagation [25], and ($ii$) the spontaneous self-focusing displayed by a beam when a particular phase distribution is imprinted in it initially [26]. In the first case, the trajectories have allowed us to provide a quantitative measure to understand how Airy-type beams degrade when we introduce conditions that diminish their ideal initial shape. Thus, we find that, the translation symmetry along the hyperbolic Bohmian trajectories disappears gradually, because there is an effective transfer of flux towards the rear part of the beam, which eventually translates also into a loss of its transverse self-acceleration. In other words, what keeps the Airy beam properties is the presence of an infinite tail. This generates a constant pressure forwards [56, 57], which is evidenced by all trajectories, regardless of their initial position, bending over at a constant rate.